\int 2^{2x}\cdot 5^x dx
1
Expert's answer
2020-09-17T14:00:06-0400

"2^{2x}\cdot 5^x=4^x\cdot 5^x=20^x"

"\int 2^{2x}\cdot 5^xdx=\int 20^xdx=\int e^{x\ln 20}dx="

"=\frac{1}{\ln 20}\int e^{x\ln 20}d(x\ln 20)=\frac{1}{\ln 20}e^{x\ln 20}+C="

"=\frac{1}{\ln 20}20^x+C"

Answer: "\frac{1}{\ln 20}20^x+C"